Dandy Manufacturers company Ltd makes of product Y has prepared it's budget for 2018, based on two activity levels of 80% and 100% with production Units of 2,800 and 3,500 units respectively.

The budget is as follows:
                                                    80%                      100%
Sales                                         $224,000             $280,000
Direct Material                        $84,000               $105,000
Direct Labour                          $50,400               $63,000
Production overhead            $48,800                $53,000

The company's actual result for the period is as follows:
Sales in Units                                                    3,150 Units
Sales                                                                $252,000
Direct material                                               $94,000
Direct labour                                                   $60,000
Overhead (60% of amount is fixed)           $50,000

Required:
Prepare a flexible budget for a production level of 70% and 90% and compare the flexible budget at 90% level of activity with the actual result.
Answer #1
Unit Sales Price=224000/2800 $            80.00 (280000/3500)
Direct Material per unit $30.00 ((8400/2800)
Direct Labor per unit $18.00 (63000/3500) (50400/2800)
FIXED PRODUCTION OVERHEAD $31,800 (53000*60%)
Variable Production overhead at100% $21,200 (53000-31800)
Variable Overhead per unit $6.06 (21200/3500)
ACTIVITY LEVEL
BUDGET 100% 80%
A Sales in units 3500 2800
B Unit Sales Price $            80.00 $             80.00
C=A*B Sales Revenue $        280,000 $         224,000
D Direct Material $105,000 $84,000
E Direct Labor $63,000 $50,400
F Fixed Production Overhead $31,800 $31,800
G Variable Overhead $21,200 $17,000 (48800-31000)
H=F+G Total Production Overhead $53,000 $48,800
I=D+E+H Total Cost $221,000 $183,200
FLEXIBLE BUDGET
ACTIVITY LEVEL
70% 90% ACTUAL Actual -Budgeted
A Sales in units 2450 3150 3150
B Unit Sales Price $            80.00 $             80.00 $        80.00
C=A*B Sales Revenue $        196,000 $         252,000 $252,000 $0
D=A*30 Direct Material $73,500 $94,500 $94,000 ($500) Favorable
E=A*18 Direct Labor $44,100 $56,700 $60,000 $3,300 Unfavorable
F=H*60% Fixed Production Overhead $31,800 $31,800 $30,000 ($1,800) Favorable
G=A*6.06 Variable Overhead $14,840 $19,080 $20,000 $920 Unfavorable
H=F+G Total Overhead $46,640 $50,880 $50,000 ($880) Favorable
I=D+E+H Total Cost $164,240 $202,080 $204,000 $1,920 Unfavorable